OpenStudy (anonymous):

write an equation in slope-intercept form of the line whose y-intercept is 4 and the line is parallel line whose equation is 6+2y=5 what is the slope intercept form? i need step by step help please

OpenStudy (anonymous):

are you sure the matching equation is 6+2y=5?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

sorry 6x+2y=5

OpenStudy (anonymous):

and y intercept is 4

OpenStudy (anonymous):

thats much better Okay so first you want to put it into slope intercept form which is y=mx+b

OpenStudy (anonymous):

okay so 2y=-6x+5 correct?

OpenStudy (anonymous):

start by subtracting 6x from both sides as it is a positive value which gives you

OpenStudy (anonymous):

2y=5-6x or 2y=-6x+5 then you divide both sides by 2 to isolate the y y=(-6x+5)/2 or y=-3x+(5/2)

OpenStudy (anonymous):

then i would get y=-3x+4 correct?

OpenStudy (anonymous):

becuse of the y intercept?

OpenStudy (anonymous):

correct because b is your 7 intercept and a parallel line has the same slope

OpenStudy (anonymous):

y intercept not 7

OpenStudy (anonymous):

your slope is m or (-3)

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

no problem